OpenStudy (anonymous):

A theater company charges $10.50 per ticket for people of all age groups. If c represents the total number of children who bought tickets and a represents the total number of adults who bought tickets, which expression best represents the total money earned by the company from the sale of tickets? 10.5c + a c + 10.5a 10.5(c + a) 21(c + a)

OpenStudy (anonymous):

@princeharryyy

OpenStudy (anonymous):

lol you were already here!

OpenStudy (princeharryyy):

:)

OpenStudy (princeharryyy):

10.5*c + 10.5*a

OpenStudy (princeharryyy):

C it is.

OpenStudy (anonymous):

well yes that should be the answer butt I don tsee it

OpenStudy (princeharryyy):

10.5(c + a)

OpenStudy (princeharryyy):

u dere ? @Blosson505

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry

OpenStudy (princeharryyy):

it's done! what's next?

OpenStudy (anonymous):

Henry wants to laminate the tabletop in his study room. The tabletop is the shape of an isosceles trapezoid shown here. If laminating costs $4 per square foot, what is the total cost to laminate the tabletop? A trapezoid with a height of 3 feet and bases of 6 feet and 10 feet is shown. $24 $72 $84 $96

OpenStudy (princeharryyy):

area of trapezoid = 1/2 * (6+10)*3 = 6*3 =24sq units money required 24*4 = $96

OpenStudy (princeharryyy):

@Blosson505

OpenStudy (princeharryyy):

what's next ?

OpenStudy (anonymous):

How is the expression (5 + y) ⋅ 8 described by its factors and terms? It has two factors and the factor (5 + y) has one term. It has three factors and the factor (5 + y) has one term. It has three factors and the factor (5 + y) has two terms. It has two factors and the factor (5 + y) has two terms.

OpenStudy (princeharryyy):

8*3 not 6*3 my bad

OpenStudy (princeharryyy):

is the expression correct

OpenStudy (anonymous):

so $96 is not the answer?

OpenStudy (princeharryyy):

answer is 96 I wrote 6 instead of 8 by mistake

OpenStudy (anonymous):

oh okay and the expression is neither right nor wrong its asking about the factors and terms

OpenStudy (princeharryyy):

is the expression correct

OpenStudy (princeharryyy):

can you printscreen the question and attach the file

OpenStudy (anonymous):

sorry I cant give me a second

OpenStudy (anonymous):

How is the expression (5 + y) ⋅ 8 described by its factors and terms? It has two factors and the factor (5 + y) has one term. It has three factors and the factor (5 + y) has one term. It has three factors and the factor (5 + y) has two terms. It has two factors and the factor (5 + y) has two terms.

OpenStudy (anonymous):

thats the question in full.....

OpenStudy (princeharryyy):

ok

OpenStudy (princeharryyy):

It has two factors and the factor (5 + y) has two terms.

OpenStudy (princeharryyy):

wait wait wait

OpenStudy (anonymous):

so teh last one?

OpenStudy (princeharryyy):

it's correct

OpenStudy (anonymous):

lol

OpenStudy (princeharryyy):

yes

OpenStudy (anonymous):

D is correct?

OpenStudy (princeharryyy):

what's next?

OpenStudy (princeharryyy):

Quickly! it's 5:07 am mommy just woke up!

OpenStudy (princeharryyy):

She has started to scold. She's gonna make me sleep. So, be fast.

OpenStudy (anonymous):

u gotta give me a sec

OpenStudy (anonymous):

A bus pass costs $8 per week. Which of the following equations shows the total cost in dollars, t, of the bus pass for a certain number of weeks, w? t = 8 + w w = 8 + t t = 8w w = 8t

OpenStudy (anonymous):

I think its C

OpenStudy (princeharryyy):

Correct! you are.

OpenStudy (princeharryyy):

next?

OpenStudy (anonymous):

Paul’s company rents cars to people who are at least 25 years old. Which graph best represents the possible ages of people who can rent cars from Paul’s company? A number line with an open circle on 25, with shading to the right. A number line with a closed circle on 25, with shading to the right. A number line with an open circle on 25, with shading to the left. A number line with a closed circle on 25, with shading to the left.

OpenStudy (princeharryyy):

where's the graph

OpenStudy (princeharryyy):

where's the graph

OpenStudy (princeharryyy):

where's the graph

OpenStudy (princeharryyy):



OpenStudy (princeharryyy):

is there any option like this

OpenStudy (princeharryyy):

I mean graph